OpenStudy (anonymous):

Solve the equation by completing the square. x^2 + 10x + 21 = 0 x= (smaller value) x= (larger value)

OpenStudy (saifoo.khan):

smaller, -7 larger -3

OpenStudy (anonymous):

x^2+10x+21=x^2+10x+25-4=0 (x+5)^2-4=0 (x+5)^2=4 So x+5=2 implies x=-3 or -(x+5)=2 implies x=-7